OpenStudy (anonymous):

solve using perfect square factoring: x^2 – 10x + 25 = 9

OpenStudy (xanthe):

Bring the 9 over

OpenStudy (anonymous):

answer a)-8, -2 b)-4, 12 c)4, -12 d)2,8

OpenStudy (nurali):

x^2 -10x +25 = 9 We're told it forms a perfect square and the first number that jumps to my mind is 5, since 5+5 is 10 and 5x5 is 25, but we still need to get the sign right. The quadratic equation says -10x, so we need the sign to be -5. -5+-5 = -10 -5*-5 = +25 So this perfect square is expressed as follows: (x-5)^2 = 9 Now we want the values for x such that (x-5)^2 = 9. (x-5) = square root of 9, wihch is 3 or -3. x = 5-3 or 5+3 x= 2 or x = 8

OpenStudy (xanthe):

Use formula